OpenStudy (anonymous):

Solve the differential equation y"-y-6y=0. For this problem, is it y"-7y=0 and you factor?

OpenStudy (anonymous):

(y-3)(y+2)=0

OpenStudy (anonymous):

are you sure you copied the problem down correctly? y"-y'-6y=0

OpenStudy (anonymous):

is this a calculus q?

OpenStudy (anonymous):

So the equation that you provided is the right problem, right? Maybe the book was wrong then.

OpenStudy (anonymous):

@Idealist eh...

OpenStudy (anonymous):

What happened?

OpenStudy (anonymous):

@kittyxie1 this is differential equations...

OpenStudy (anonymous):

If it's \(y''-y-6y=0\) then yes it's identical to \(y''-7y=0\) but why would they give it unsimplified?

OpenStudy (anonymous):

make the substitution \[y= e^{\lambda t}\]

OpenStudy (anonymous):

y"=7y, is this double derivative?

OpenStudy (anonymous):

y'' is the 2nd derivative of y

OpenStudy (anonymous):

so yes

OpenStudy (anonymous):

\[y = e^{\lambda t} \] \[y'' =\lambda ^2 e^{\lambda t}\]

OpenStudy (anonymous):

it's easier to recognize this as an eigenvalue problem

OpenStudy (anonymous):

y'=7y^2/2 y=7y^3/6

OpenStudy (anonymous):

y=7y^3/6+c

OpenStudy (anonymous):

@kittyxie1 your answers make no sense, if you dont know how to solve differential equations please dont attempt to do so

OpenStudy (anonymous):

\[y'' - 7 y = 0\] \[y= e^{ \lambda t}\] substituting for y \[\lambda ^2 e^{ \ \lambda t} - 7e^{\lambda t} = 0\] factoring out the exponent \[(\lambda ^2 -7) e ^ {\lambda t}=0\] now its impossible for the exponent to be zero which means \[\lambda ^2 -7 = 0\] or you can get to this point by dividing the exponents to both sides solve for lambda then write it in the form of \[y= Ae^{\lambda _ 1 t}+Be^{\lambda _2 t}\] if there are 2 different solutions for lambda

OpenStudy (anonymous):

@kittyxie1 fyi

OpenStudy (anonymous):

I want to thank you guys so much for helping me. Thanks a lot, guys.

OpenStudy (anonymous):

if there is a single root when solving out for lambda it will look like \[y= Ae^{\lambda t} + B t e^{\lambda t}\] dont ask what if the solution is complex because i dont remember the form off the top of my head

OpenStudy (anonymous):

@completeidiot that's rude... you needn't attack @kittyxie1 for being wrong

OpenStudy (anonymous):

@oldrin.bataku its not meant to be an attack, last thing we want to do is mislead the asker

OpenStudy (anonymous):

btw the 'form' is actually pretty simple... if we have \(\lambda=a+bi\):$$y=Ae^{(a+bi)t}=Ae^ae^{bit}=(Ae^a)e^{bit}=(Ae^a)(\cos bt+i\sin bt)=A_1\cos bt+A_2\sin bt$$

OpenStudy (anonymous):

yeah that totally makes sense